OpenStudy (anonymous):

12=-7+y solve and check with full workins cos my proffesor requested for full workings

OpenStudy (anonymous):

Add 7 both sides, done.

OpenStudy (anonymous):

wt bt check

OpenStudy (anonymous):

What u want me to check?

OpenStudy (anonymous):

for the eqn 12=-7+y

OpenStudy (anonymous):

What about it?

OpenStudy (anonymous):

Still don't understand what it is u want me to check..?

OpenStudy (anonymous):

-5+t=11 solve and check

OpenStudy (anonymous):

Add 5 both sides, what's to check?

OpenStudy (anonymous):

I saw in the other questions that u want me to add 5, then check that I added 5 correctly by taking it away again....which is pretty stupid, so I will leave the adding and subtracting to u if that's OK.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

r+1/3=8/3-1/3 solve and check

OpenStudy (anonymous):

This just same thing, take away 1/3 both sides...

OpenStudy (anonymous):

so how do u check if d eqn is right

OpenStudy (anonymous):

Look, take your example -5+t=11 solve and check If I was doing this myself, I would just write t=16 (full stop). Since u requested working, I said "add 5 to each side" . If u do that correctly, then the original equation is not changed. If u want to check that u added 5 correctly, do it again as many times as it takes to make you feel certain that u added 5 correctly,

OpenStudy (anonymous):

r u mad@ me.ok im sorry 4 asking u question ok

OpenStudy (anonymous):

u might as well ignore my question and let others reply.once again im sorry

OpenStudy (anonymous):

I am not mad at u. U have asked me several times now to check something when I haven't even done anything, I asked u to do it!!!!!

OpenStudy (anonymous):

what u need help with? :)

OpenStudy (anonymous):

r+1/3=8/3-1/3 r = 7/3 - 1/3 r = 6/3 r = 1/2 to check, plug 1/2 in for r and see if that side of the equation equals the other side (8/3 - 1/3)